If an ISP has a 10.0.0.0/8 you are required to give 125, 30, 500, 6, 2, 2, 2 and host addresses. Supply the needed information in table given.Solve and show solution.

Hosts n   

Prefix Length(32-n)

Max of hosts(2n – 2)

Inc 2n

Network IP Subnet Mask Broadcast IP
Answer #1

Arrange the hosts requirements in descending order and assign the highest IP to the highest host requirement.

500

125

30

6

2

2

2

For 500 hosts addresses, 9(2^9=512) bits are required for the host part to uniquely identify each host. The remaining 32-9-8=15 bits are used for subnet part. Thus the subnet mask is /23(8+15).

For 125 hosts addresses, 7(2^7=126) bits are required for host part to uniquely identify each host. The remaining 32-7-8=17 bits are used for subnet part. Thus the subnet mask is/25(8+17).

For 30 hosts addresses, 5(2^5=32) bits are required for host part to uniquely identify each host. The remaining 32-5-8=19 bits are used for subnet part. Thus the subnet mask is /27(8+19).

For 6 hosts addresses, 3(2^3=8) bits are required for host part to uniquely identify each host. The remaining 32-3-8=21 bits are used for subnet part. Thus the subnet mask is /29(8+21).

For 2 hosts addresses, 2(2^2=4) bits are required for host part to uniquely identify each host. The remaining 32-2-8=22 bits are used for subnet part. Thus the subnet mask is/30(8+22).

Hosts n Prefix Length Max of hosts Inc 2n Network IP Subnet Mask Broadcast
500 9 32-9=23 512-2=510 512 10.0.0.0 /23 10.0.1.255
125 7 32-7=25 128-2=126 128 10.0.2.0 /25 10.0.2.127
30 5 32-5=27 32-2=30 32 10.0.2.128 /27 10.0.2 159
6 3 32-3=29 8-2=6 8 10.0.2.160 /29 10.0.2.167
2 2 32-2=30 4-2=2 4 10.0.2.168 /30 10.0.2.171
2 2 32-2=30 4-2=2 4 10.0.2.172 /30 10.0.2.175
2 2 32-2=30 4-2=2 4 10.0.2.176 /30 10.0.2.179